OpenStudy (anonymous):

What is the equation, in standard form, of a vertical ellipse centered at (–8, 12) with a major axis of length 10 and a minor axis of length…

OpenStudy (dumbcow):

minor axis of length what?? that would be needed to complete equation

OpenStudy (dumbcow):

\[\frac{(y-k)^{2}}{a^{2}}+\frac{(x-h)^{2}}{b^{2}} = 1\] center (h,k) length major axis = 2a minor axis = 2b